The width of a rectangle is 5 inches less than the length. The perimeter is 38 inches. Find the length and the width.

Answer by fractalier(6550)

Perimeter P = 2L + 2W = 38. We also have
W = L - 5
Substitute this into the first equation and get
2L + 2(L - 5) = 38
2L + 2L - 10 = 38
4L - 10 = 38
4L = 48
L = 12 inches
W = 7 inches